$9000 loan is to be repaid over 7 years with monthly payments of J12=12% p.a. Determine the amount of principal that is paid on the 5 payment.
1
Expert's answer
2020-06-23T19:43:31-0400

Monthly payment = p (r(1+r)n)/((1+r)n-1)


Mp =9000(0.01(1+0.01)84)/((1+0.01)83)


Mp=9000(0.01(1.01)84)/(1.01)83)


Mp= 9000(0.023)/(2.284)


Mp= 90.63


1St Payment:


9000-90.63=8909.37


2Nd Payment:


8909.37-9063= 8818.74


3rd payment:


8818.74-90.63= 8728.11


4th payment:

8728.11-90.63= 8637.48


5th payment:


8637.48-90.63= 8546.85